A 2kg box is launched up a ramp with sufficient speed to reach the top. The ramp is 3m long, and inclined at 20dg. The coefficient of kinetic friction between the package and the plane is 0.4. What is the minimum kinetic energy required for the package?
Answer #1

Date 0 = 20° 2 kg l=3m Mp = 0.4 Drawing FBD of box fx اسباني ngcose resnio perpendicular to ramp. Ef=o (no motion) Namgcoco =Now

K.E. = \frac{1}{2}mu^2

K.E. = \frac{1}{2}*2*6.5^2

K.E. =42.25 J